# Multi-Grained Query-Guided Set Prediction Network for Grounded Multimodal Named Entity Recognition

### v2

@@ Abstract
Grounded Multimodal Named Entity Recognition (GMNER) is an emerging information extraction (IE) task, aiming to simultaneously extract entity spans, types, and corresponding visual regions of entities from given sentence-image pairs data. Recent unified methods employing machine reading comprehension or sequence generation-based frameworks show limitations in this difficult task. The former, utilizing human-designed queries, struggles to differentiate ambiguous entities, such as Jordan (Person) and off-White x Jordan (Shoes). The latter, following the one-by-one decoding order, suffers from exposure bias issues. We maintain that these works misunderstand the relationships of multimodal entities. To tackle these, we propose a novel unified framework named Multi-grained Query-guided Set Prediction Network (MQSPN) to learn appropriate relationships at intra-entity and inter-entity levels. Specifically, MQSPN consists of a Multi-grained Query Set (MQS) and a Multimodal Set Prediction Network (MSP). MQS explicitly aligns entity regions with entity spans by employing a set of learnable queries to strengthen intra-entity connections. Based on distinct intra-entity modeling, MSP reformulates GMNER as a set prediction, guiding models to establish appropriate inter-entity relationships from a global matching perspective. Additionally, we incorporate a query-guided Fusion Net (QFNet) to work as a glue network between MQS and MSP. Extensive experiments demonstrate that our approach achieves state-of-the-art performances in widely used benchmarks [^1] Code is available at https://github.com/tangjielong928/mqspn..
@@ Introduction
To effectively comprehend and manage vast amounts of multimodal content from social media, recent research (Yu et al. 2023) proposes a nascent multimodal information extraction (IE) task named Grounded Multimodal Named Entity Recognition (GMNER), which aims to extract multimodal entity information, including entity spans, types, and corresponding regions of entities, from image-text pairs. Prior studies decompose the GMNER task into several subtasks such as Multimodal Named Entity Recognition (Zhang et al. 2021a) and Visual Grounding (Redmon and Farhadi 2018), adopting the pipeline approach to solve it, which leads to serious error propagation. To address this problem, recent research paradigm has transferred to detect span-type-region triplets with unified model by formulating GMNER as machine reading comprehension (MRC) (Jia et al. 2023) or sequence generation (Yu et al. 2023; Wang et al. 2023).
Despite their remarkable performance, limitations still exist. The MRC-based frameworks utilize human-designed queries as prior instructions to simultaneously guide entity recognition and entity grounding, struggling to distinguish different ambiguous entities. For example in Figure 1 (a), with the input of multiple fixed person queries like “Please extract person: People’s name…”, the model incorrectly detects off-White x Jordan (Shoes) as Jordan (Person) and assigns the wrong region originally belonging to Kevin Durant. On the other hand, sequence generation-based methods suffer from exposure bias. They autoregressively decode span-type-region triples one by one in predefined sequence order, resulting in the prediction of off-White x Jordan region highly sensitive to errors in preceding Kevin Durant detection in Figure 1 (b).
@@
In our perspective, the essence of these errors is the inappropriate modeling of multimodal entity relationships. (1) MRC-based frameworks overlook the distinctions of intra-entity relationships due to their reliance on duplicate and untrainable queries that cannot distinguish intrinsic features of ambiguous entities. (2) Sequence generation-based methods excessively rely on inter-entity relationships between different multimodal entities, where the current output is vulnerable to previous predictions. Therefore, we propose a novel unified framework named MQSPN, mainly consisting of learnable Multi-grained Query Set (MQS) and Multimodal Set Prediction Network (MSP) to focus on modeling appropriate relationships at intra-entity and inter-entity levels.
We maintain that modeling intra-entity connections is foundational; only when a model is capable of distinguishing individual entities can it further model inter-entity relationships effectively. Hence, we propose MQS to adaptively learn intra-entity connections. Instead of human-designed queries in the MRC-based framework, MQS adopts a set of learnable queries (Li et al. 2023; Gong et al. 2024) (denoted as entity-grained query) to perform joint learning in span localization, region grounding, and entity classification for different entities, which enforces queries to learn distinguishable features and automatically establish explicit intra-entity connections. However, simply learnable queries are insufficient to detect regions and spans due to the lack of semantics information. To tackle this, we feed a prompt with masked type statement into vanilla BERT to build type-grained queries with type-specific semantics. Finally, each multi-grained query is constructed by integrating a learnable entity-grained query with a type-grained query.
@@
Besides, since directly fusing textual features with irrelevant visual features will impair model performance (Chen et al. 2021), we further propose a QFNet between MQS and MSP to filter this noisy information, thereby boosting better alignment of two-level relationships. Unlike direct fusion methods (Yu et al. 2020; Zhang et al. 2021b; Wu et al. 2023), QFNet employs queries as intermediaries to facilitate the separate integration of textual and visual region representations. More details are illustrated in the following section. Our contributions could be summarized as follows:
@@
Grounded Multimodal Named Entity Recognition. Previous multimodal named entity recognition (MNER) models (Yu et al. 2020; Zhang et al. 2021b; Chen et al. 2022a; Zhang et al. 2021a; Wang et al. 2022b) primarily focused on how to utilize visual information to assist textual models in entity extraction. Grounded MNER (GMNER) (Yu et al. 2023) is proposed to additionally output the bounding box coordinates of named entities within the image. The taxonomy of previous GMNER works encompasses two branches, namely pipeline manner and unified manner. Pipeline methods (Li et al. 2024a; Ok et al. 2024) decompose the GMNER task into several subtasks like MNER, entity entailment and visual grounding (Wang et al. 2022a). To tackle the error propagation issue, unified methods formulate GMNER as an end-to-end machine reading comprehension (Jia et al. 2023, 2022) or sequence generation task (Wang et al. 2023; Yu et al. 2023; Li et al. 2024b). Different from them, our MQSPN reformulate GMNER as set prediction to learn appropriate intra-entity and inter-entity relationships.
Set Prediction. Set prediction is a well-known machine learning technique where the goal is to predict an unordered set of elements. It is widely used in various applications such as object detection (Carion et al. 2020), point cloud classification (Qi et al. 2017), and multi-label classification (Yang et al. 2019). In the field of NER, Seq2Set (Tan et al. 2021) first reformulates the nested NER task as set prediction to eliminate order bias from the autoregressive model. Subsequently, PIQN (Shen et al. 2022) associates each entity with a learnable instance query to capture the location and type semantics. Recently, DQSetGen (Chen et al. 2024) extended set prediction to other sequence labeling tasks such as slot filling and part-of-speech (POS) Tagging. Different from these methods, we introduce the MQSPN, which exploits the set prediction paradigm in a new GMNER area to model the two-level relationship of multimodal entities.
Figure 2: (a). Overview of our MQSPN. MQS and MSP guide models to learn intra-entity and inter-entity relationships respectively. (b). The construction of Multi-grained Query Set (MQS) consists of Type-grained Queries and Learnable Entity-grained Queries (LEQ). (c). The detailed architecture of Query-guided Fusion Net (QFNet).
@@
Visual Representation. Given the input image $I$, we utilize VinVL (Zhang et al. 2021c) as the class-agnostic region proposal network (RPN) to obtain all candidate regions. Following the work of (Yu et al. 2023), we also retain the top-$k$ region proposals as our candidate regions, using the ViT-B/32 (Dosovitskiy et al. 2020) from pre-training CLIP (Radford et al. 2021) as vision encoder. The initial visual representation for the candidate regions are denoted as $\mathbf{V}=\{\mathbf{v}_{1},\ldots,\mathbf{v}_{k}\}$. To match those entities that are ungroundable in the images, we construct a special visual token embedding $\mathbf{v}_{[ug]}$ by feeding a blank image $\mathcal{T}$ into vision encoder. Finally, $\mathbf{v}_{[ug]}$ and $\mathbf{V}$ are concatenated to serve as the final visual representation $H_{V}\in\mathbb{R}^{(k+1)\times h}$, where $k$ is the number of candidate regions.
@@
Table 1: Performance comparison of different competitive approaches on Twitter-GMNER datasets. Bold represents the optimal result, and underlined represents the suboptimal result. For the baseline methods, $\clubsuit$ indicates that the results are reproduced according to the corresponding papers, and others are from (Yu et al. 2023).
@@
GMNER Datasets. We conduct experiments on the Twitter-GMNER (Yu et al. 2023) and Twitter-FMNERG (Wang et al. 2023). Please refer to Appendix for detailed information about these datasets.
Evaluation Metrics. Following previous works (Yu et al. 2023), we use Precision (Pre.), Recall (Rec.), and F1 Score (F1) as the evaluation metric for GMNER and its two subtasks. See Appendix for a more detailed definition.
Implementation Details All experiments are implemented on 4 NVIDIA RTX3090 GPUs with Pytorch 1.9.1. For a fair comparison, we use the pre-trained BERT-based model [^2] https://huggingface.co/google-bert/bert-base-uncased as the textual encoder, ViT-B/32 from pre-training CLIP [^3] https://huggingface.co/openai/clip-vit-base-patch32 as the visual encoder, and VinVL [^4] https://github.com/pzzhang/VinVL as a class-agnostic RPN. For model’s efficiency, we freeze the visual encoder and RPN and assign the layer of the QFNet as $L=3$. We initialize the learnable query part with a normal distribution $\mathcal{N}(0,0.02)$. For training, we set the batch size to 16, the learning rate to $2\times 10^{-5}$, and the training epoch to 50. Our model uses an Adam optimizer with a linear warmup of ratio 0.05. We tune the optimal query number $u$ to 60 and 120 for Twitter-GMNER and Twitter-FMNERG datasets, respectively. To allow the multi-grained queries to learn query semantics initially, we first train the model for 5 epochs with the freezing parameters of the pre-training model. Please refer to Appendix for baseline methods.
Table 2: F1 scores of previous SOTA methods H-Index, TIGER, and our MQSPN on Twitter-FMNERG datasets for each fine-grained entity type in GMNER task.
Entity Type H-Index TIGER MQSPN Person 45.13 43.78 48.64 Location 62.33 67.69 71.03 Building 32.88 40.00 44.78 Organization 46.68 46.75 49.03 Product 28.19 27.38 29.46 Art 38.89 43.27 42.84 Event 45.56 48.39 51.88 Other 41.81 48.28 43.49 All 46.55 47.20 48.57
@@
Performance on GMNER, MNER, and EEG. Following (Yu et al. 2023), we also report two subtasks of GMNER, i.e., Multimodal Named Entity Recognition (MNER) and Entity Extraction & Grounding (EEG). MNER aims to identify entity span and entity type pairs, while EEG aims to extract entity span and entity region pairs. Table 1 shows the performance comparison of our method with competitive baseline models on Twitter-GMNER benchmarks.
@@
Table 3: Ablation study of each component on overall F1 score of different tasks.
Module Settings Twitter-GMNER Twitter-FMNERG GMNER MNER EEG GMNER MNER EEG MQSPN Default 58.76 80.43 62.40 48.57 67.09 62.64 MQS w/o PTQ 56.95 ($\downarrow$1.81) 79.78 61.51 47.23 ($\downarrow$1.34) 65.86 61.75 w/o LEQ 56.02 ($\downarrow$2.74) 78.64 60.30 45.39($\downarrow$3.18) 62.91 59.39 QFNet w/o QCT 57.21 ($\downarrow$1.55) 80.04 61.88 47.61($\downarrow$0.96) 66.14 62.24 w/o QPI 57.86 ($\downarrow$0.90) 80.11 61.91 48.33($\downarrow$0.24) 66.59 62.55 w/o SAG 58.13 ($\downarrow$0.63) 80.23 62.26 47.96($\downarrow$0.61) 66.37 62.37 MSP w/o BML 56.58 ($\downarrow$2.18) 79.26 61.37 46.81($\downarrow$1.76) 64.43 60.81
@@
Analysis of query quantity. To explore the impact of multi-grained query quantity (i.e., $u$) on GMNER and its 2 subtasks, we report the F1 scores of Twitter-GMNER in Figure 4 by tuning $u$ from $10$ to $90$. We observe that as the number of queries increases from $10$ to $60$, the model’s performance improves by $+1.90\%$, $+1.76\%$, and $+1.45\%$ in the GMNER, MNER, and EEG tasks, respectively. These results indicate that query quantity plays a crucial role in intra-entity connection learning. However, a query quantity exceeding $60$ does not lead to better performance. This suggests that there is an optimal number of queries for MQSPN. In our experiments, we set $u=60$ in the Twitter-GMNER dataset to achieve the optimal results.
Figure 5: (a). Predictions of MQSPN in different ablation setting, which is the same as Table 3. (b). Predictions of TIGER and our MQSPN. The middle and right parts are the visualizations of attention weights in MQSPN. The x-axis denotes the MQS and the y-axis represents sentence tokens or candidate visual regions.
Case Study. We conduct a comprehensive case study in two samples. As shown in Figure 7 (a), when we replace learnable queries with human-designed queries in MQS (w/o LEQ setting), the model cannot differentiate Bradon Saad and Johan Franzen regions, and incorrectly localize Blackhawks and Detroit as Person spans, highlighting deficiencies in intra-entity learning for ambiguous entities. After integrating learnable MQS, the model precisely aligns textual spans with corresponding regions for Person entities. However, using MQS without MSP (w/o BML) leads to confusion in mapping proper inter-entity relationships, resulting in incorrect classifications for Blackhawks and Detroit and more erroneous span-type-region triplets. After combining MQS and MSP, MQSPN effectively differentiates entities and makes accurate predictions, demonstrating the effect of intra-entity and inter-entity modeling.
In Figure 7 (b), TIGER incorrectly identifies the ambiguous entity Jimmy Kimmel within Jimmy Kimmel Live as Person, and associates it with the error region. This results in incorrectly matching the Region-2 with The Weeknd, which originally belongs to Region-0. In contrast, our MQSPN can establish correct intra-entity connections between all textual spans and regions. To further investigate the contribution of MQS, we perform visualizations of the query-text and query-regions attention maps. We find that type-specific queries have the highest attention weights in their target spans or regions, which indicates that type-grained queries of MQS provide effective information. Meanwhile, we observe that different queries are responsible for identifying different spans and regions. For instance, the first PER (person) query mainly focuses on Donald Trump and Region-1, while the second one has the highest attention scores on The Weeknd and Region-0. This demonstrates that these trainable entity-grained queries of MQS adaptively learn distinguishable features for different entities to establish more stable intra-entity connections.
More experiments are attached in the Appendix.

### v3

@@ Abstract
Grounded Multimodal Named Entity Recognition (GMNER) is an emerging information extraction (IE) task, aiming to simultaneously extract entity spans, types, and corresponding visual regions of entities from given sentence-image pairs data. Recent unified methods employing machine reading comprehension or sequence generation-based frameworks show limitations in this difficult task. The former, utilizing human-designed type queries, struggles to differentiate ambiguous entities, such as Jordan (Person) and off-White x Jordan (Shoes). The latter, following the one-by-one decoding order, suffers from exposure bias issues. We maintain that these works misunderstand the relationships of multimodal entities. To tackle these, we propose a novel unified framework named Multi-grained Query-guided Set Prediction Network (MQSPN) to learn appropriate relationships at intra-entity and inter-entity levels. Specifically, MQSPN explicitly aligns textual entities with visual regions by employing a set of learnable queries to strengthen intra-entity connections. Based on distinct intra-entity modeling, MQSPN reformulates GMNER as a set prediction, guiding models to establish appropriate inter-entity relationships from a optimal global matching perspective. Additionally, we incorporate a query-guided Fusion Net (QFNet) as a glue network to boost better alignment of two-level relationships. Extensive experiments demonstrate that our approach achieves state-of-the-art performances in widely used benchmarks.
Codehttps://github.com/tangjielong928/mqspn
@@ Introduction
To effectively comprehend and manage vast amounts of multimodal content from social media, recent research (Yu et al. 2023a) proposes a nascent multimodal information extraction task named Grounded Multimodal Named Entity Recognition (GMNER). It aims to extract multimodal entity information, including entity spans, types, and corresponding regions of entities, from image-text pairs. Prior studies decompose the GMNER task into several subtasks such as Multimodal Named Entity Recognition (Zhang et al. 2021a) and Visual Grounding (Redmon and Farhadi 2018), adopting the pipeline approach to solve it, which leads to serious error propagation. To address this problem, recent research paradigm has transferred to detect span-type-region triplets with unified model by formulating GMNER as machine reading comprehension (MRC) (Jia et al. 2023) or sequence generation (Yu et al. 2023a; Wang et al. 2023).
Despite their remarkable performance, limitations still exist. The MRC-based frameworks utilize human-designed type-specific queries as prior instructions to simultaneously guide entity recognition and entity grounding, struggling to distinguish different ambiguous entities. For example in Figure 1 (a), with the input of multiple fixed person queries like ”Please extract person: People’s name…”, the model incorrectly detects off-White x Jordan (Shoes) as Jordan (Person) and assigns the wrong region originally belonging to other entity (Kevin Durant). On the other hand, sequence generation-based methods suffer from exposure bias. They autoregressively decode span-type-region triples one by one in predefined sequence order, resulting in the prediction of off-White x Jordan region highly sensitive to errors in preceding Kevin Durant detection in Figure 1 (b).
@@
In our perspective, the essence of these errors is the inappropriate modeling of multimodal entity relationships. (1) Manually building specific query for each entity is labor-intensive and impractical. Existing MRC-based frameworks overlook the distinctions of intra-entity connection due to their reliance on duplicate and untrainable type queries that cannot distinguish intrinsic features of ambiguous entities. (2) Sequence generation-based methods excessively rely on inter-entity relationships between different multimodal entities elements, where the current output is vulnerable to previous predictions. Therefore, we propose a novel unified framework named MQSPN, mainly consisting of learnable Multi-grained Query Set (MQS) and Multimodal Set Prediction Network (MSP) to focus on modeling appropriate relationships at intra-entity and inter-entity levels.
We maintain that modeling intra-entity connections is foundational; only when a model is capable of distinguishing individual entities can it further model inter-entity relationships effectively. Hence, we first propose MQS to adaptively learn intra-entity connections. Instead of human-designed queries in the MRC-based framework, MQS adopts a set of learnable queries (Li et al. 2023; Gong et al. 2024) (denoted as entity-grained query) to perform joint learning in span localization, region grounding, and entity classification for different entities, which enforces queries to learn distinguishable features and automatically establish explicit intra-entity connections. However, simply learnable queries are insufficient to detect regions and spans due to the lack of semantics information. To tackle this, we feed a prompt with masked type statement into vanilla BERT to build type-grained queries with type-specific semantics. Finally, each multi-grained query is constructed by integrating a learnable entity-grained query with a type-grained query.
@@
Besides, since directly fusing textual features with irrelevant visual features will impair model performance (Chen et al. 2021), we further propose a QFNet between MQS and MSP to filter this noisy information, thereby boosting better alignment of two-level relationships. Unlike direct fusion methods (Zhang et al. 2021b; Wu et al. 2023), QFNet employs queries as intermediaries to facilitate the separate integration of textual and visual region representations. Our contributions could be summarized as follows:
@@
Grounded Multimodal Named Entity Recognition. Previous multimodal named entity recognition (MNER) models (Yu et al. 2020; Chen et al. 2022a; Zhang et al. 2021a) primarily focused on how to utilize visual information to assist textual models in entity extraction. Grounded MNER (GMNER) (Yu et al. 2023a) is proposed to additionally output the bounding box coordinates of named entities within the image, which has great potential in various downstream tasks, such as knowledge base construction (Liu et al. 2019) and QA systems (Yu et al. 2021; Yu, Zha, and Yin 2019; Yu et al. 2023b). The taxonomy of previous GMNER works encompasses two branches, namely pipeline manner and unified manner. Pipeline methods (Li et al. 2024a; Ok et al. 2024) decompose the GMNER task into several subtasks like MNER, entity entailment and entity grounding. To tackle the error propagation issue, unified methods formulate GMNER as an end-to-end machine reading comprehension (Jia et al. 2023, 2022) or sequence generation task (Wang et al. 2023; Yu et al. 2023a; Li et al. 2024b). Different from them, our MQSPN reformulate GMNER as set prediction to learn appropriate intra-entity and inter-entity relationships.
Set Prediction. Set prediction is a well-known machine learning technique where the goal is to predict an unordered set of elements. It is widely used in various applications such as object detection (Carion et al. 2020). In the field of NER, Seq2Set (Tan et al. 2021) first reformulates the nested NER task as set prediction to eliminate order bias from the autoregressive model. Subsequently, PIQN (Shen et al. 2022) associates each entity with a learnable instance query to capture the location and type semantics. Recently, DQSetGen (Chen et al. 2024) extended set prediction to other sequence labeling tasks such as slot filling and part-of-speech tagging. Different from these methods, we introduce the MQSPN, which exploits the set prediction paradigm in a new GMNER area to model the two-level relationship of multimodal entities.
Figure 2: (a). Overview of our MQSPN. (b). The construction of Multi-grained Query Set (MQS) consists of Type-grained Queries and Learnable Entity-grained Queries (LEQ). (c). The detailed architecture of Query-guided Fusion Net (QFNet).
@@
Visual Representation. Given the input image $I$, we utilize VinVL (Zhang et al. 2021c) as the class-agnostic region proposal network (RPN) to obtain all candidate regions. Following the work of (Yu et al. 2023a), we also retain the top-$k$ region proposals as our candidate regions, using the ViT-B/32 (Dosovitskiy et al. 2020) from pre-training CLIP (Radford et al. 2021) as vision encoder. The initial visual representation for the candidate regions are denoted as $\mathbf{V}=\{\mathbf{v}_{1},\ldots,\mathbf{v}_{k}\}$. To match those entities that are ungroundable in the images, we construct a special visual token embedding $\mathbf{v}_{[ug]}$ by feeding a blank image $\mathcal{T}$ into vision encoder. Finally, $\mathbf{v}_{[ug]}$ and $\mathbf{V}$ are concatenated to serve as the final visual representation $H_{V}\in\mathbb{R}^{(k+1)\times h}$, where $k$ is the number of candidate regions.
@@
Category Methods GMNER MNER EEG
Pre. Rec. F1 Pre. Rec. F1 Pre. Rec. F1
Pipeline Methods GVATT-RCNN-EVG 49.36 47.80 48.57 78.21 74.39 76.26 54.19 52.48 53.32
UMT-RCNN-EVG 49.16 51.48 50.29 77.89 79.28 78.58 53.55 56.08 54.78
UMT-VinVL-EVG 50.15 52.52 51.31 77.89 79.28 78.58 54.35 56.91 55.60
UMGF-VinVL-EVG 51.62 51.72 51.67 79.02 78.64 78.83 55.68 55.80 55.74
ITA-VinVL-EVG 52.37 50.77 51.56 80.40 78.37 79.37 56.57 54.84 55.69
BARTMNER-VinVL-EVG 52.47 52.43 52.45 80.65  80.14 80.39 55.68 55.63 55.66
@@
H-Index 56.16 56.67 56.41 79.37 80.10 79.73 60.90 61.46 61.18
[…]
MQSPN (Ours) 59.03 58.49 58.76 81.22 79.66 80.43 61.86 62.94 62.40
Table 1: Performance comparison of different competitive approaches on Twitter-GMNER datasets. Bold represents the optimal result, and underlined represents the suboptimal result. For the baseline methods, $\clubsuit$ indicates that the results are reproduced according to the corresponding papers, and others are from (Yu et al. 2023a).
@@
Entity Type H-Index TIGER MQSPN
Person 45.13 43.78 48.64
Location 62.33 67.69 71.03
Building 32.88 40.00 44.78
Organization 46.68 46.75 49.03
Product 28.19 27.38 29.46
Art 38.89 43.27 42.84
Event 45.56 48.39 51.88
Other 41.81 48.28 43.49
All 46.55 47.20 48.57
Table 2: F1 scores of previous unified methods and our MQSPN on Twitter-FMNERG datasets for each fine-grained entity type in GMNER task.
@@
GMNER Datasets. We conduct experiments on the Twitter-GMNER (Yu et al. 2023a) and Twitter-FMNERG (Wang et al. 2023). Please refer to Appendix for detailed information about these datasets and evaluation metrics.
Implementation Details All experiments are implemented on 4 NVIDIA RTX3090 GPUs with Pytorch 1.9.1. For a fair comparison, we use the pre-trained BERT-based model [^1] https://huggingface.co/google-bert/bert-base-uncased as the textual encoder, ViT-B/32 from pre-training CLIP [^2] https://huggingface.co/openai/clip-vit-base-patch32 as the visual encoder, and VinVL [^3] https://github.com/pzzhang/VinVL as a class-agnostic RPN. For model’s efficiency, we freeze the visual encoder and RPN and assign the layer of the QFNet as $L=3$. We initialize the learnable query part with a normal distribution $\mathcal{N}(0,0.02)$. For training, we set the batch size to 16, the learning rate to $2\times 10^{-5}$, and the training epoch to 50. Our model uses an Adam optimizer with a linear warmup of ratio 0.05. To allow the multi-grained queries to learn query semantics initially, we first train the model for 5 epochs with the freezing parameters of the pre-training model. Please refer to Appendix for baseline methods.
@@
Performance on GMNER, MNER, and EEG. Following (Yu et al. 2023a), we also report two subtasks of GMNER, i.e., Multimodal Named Entity Recognition (MNER) and Entity Extraction & Grounding (EEG). MNER aims to identify entity span and entity type pairs, while EEG aims to extract entity span and entity region pairs. Table 1 shows the performance comparison of our method with competitive baseline models on Twitter-GMNER benchmarks.
@@
Module Settings GMNER FMNERG
MQSPN - 58.76 48.57
MQS w/o PTQ 56.95 ($\downarrow$1.81) 47.23 ($\downarrow$1.34)
w/o LEQ 56.02 ($\downarrow$2.74) 45.39($\downarrow$3.18)
QFNet w/o QCT 57.21 ($\downarrow$1.55) 47.61($\downarrow$0.96)
w/o QPI 57.86 ($\downarrow$0.90) 48.33($\downarrow$0.24)
w/o SAG 58.13 ($\downarrow$0.63) 47.96($\downarrow$0.61)
MSP w/o BML 56.58 ($\downarrow$2.18) 46.81($\downarrow$1.76)
Table 3: Ablation study of each component on overall F1 score of Twitter-GMNER and Twitter-FMNERG.
Figure 5: Predictions of MQSPN in different ablation setting.
@@
Analysis of query quantity. To explore the impact of multi-grained query quantity (i.e., $u$) on GMNER and its 2 subtasks, we report the F1 scores of Twitter-GMNER in Figure 4 by tuning $u$ from $10$ to $90$. We observe that as the number of queries increases from $10$ to $60$, the model’s performance gradually improves. These results indicate that query quantity plays a crucial role in intra-entity connection learning. However, a query quantity exceeding $60$ does not lead to better performance. There is an optimal number of queries for MQSPN. In our experiments, we find $u=60$ in the Twitter-GMNER dataset can achieve the optimal results.
Case Study. We conduct a comprehensive case study in different ablation setting. As shown in Figure 7, when we replace learnable queries with human-designed queries in MQS (w/o LEQ setting), the model cannot differentiate Bradon Saad and Johan Franzen regions, and incorrectly localize Blackhawks and Detroit as Person spans, highlighting deficiencies in intra-entity learning for ambiguous entities. After integrating learnable MQS, the model precisely aligns textual spans with corresponding regions for Person entities. However, using MQS without MSP (w/o BML) leads to confusion in mapping proper inter-entity relationships, resulting in incorrect classifications for Blackhawks and Detroit and more erroneous span-type-region triplet predictions. After combining MQS and MSP, MQSPN effectively differentiates entities and makes accurate predictions, demonstrating the effect of intra-entity and inter-entity modeling.